OpenStudy (anonymous):

Write an equation in standard form for a line passing through A(–2, 3) and B(3, 4).

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

first find slope using the slope formula given, then substitute the slope and one of the 2 points into the pointslope formula given then convert that into standard form

OpenStudy (anonymous):

hi stevenc did you get what complete have explain to you? or you need more explanation?

OpenStudy (anonymous):

I plugged it into point slope and got y-3= 1/5 (x-4)

OpenStudy (anonymous):

@mark_o.

OpenStudy (anonymous):

ok from y-y1=m(x-x1), if you have chosen the point of A(-2,3) it should be y-y1=m(x-x1) y-3=1/5 (x-(-2)) y-3=1/5 (x+2)

OpenStudy (anonymous):

I see what I did wront now, but now im still not getting any of the answer choices

OpenStudy (anonymous):

you will also arrive to the same answer if you have chosen the point B(3,4)=(x1,y1) y-y1=m(x-x1) y-4=1/5 (x-3)

OpenStudy (anonymous):

The possible answer choices are 5x – y = 17 x – y = 1 x – 5y = 19 x – 5y = –17

OpenStudy (anonymous):

yes i know i am just telling you how to start solving it, let say you chose A(-2,3) then from the above solution we get y-3=1/5 (x+2) doing the process of multiplying them and adding y-3=1/5 x+2/5 y=1/5 x+2/5 +3 y=1/5 x+2/5 +15/5 now we want the form of this Ax+By=C y=1/5 x +17/5 now multiply both sides by 5 we get 5y=x+17 arrange them like this std form Ax+By=C -17=x-5y or x-5y=-17 .....ans....

OpenStudy (anonymous):

is that solution help u a lot now?

OpenStudy (anonymous):

oh okay!! Thanks! I didn't do the multiply both sides by 5, but I understant now. Thanks again!

OpenStudy (anonymous):

ok yw have fun .....lol